OpenStudy (anonymous):

-2/5 - 7/10 = ? (Show Work)

hero (hero):

Hint: Multiply the first fraction by 2/2

OpenStudy (anonymous):

-2/5 - 7/10 = x *2 -4/10-7/10=x Answere is -11/10

OpenStudy (anonymous):

same as the last. make the denominators equal. -2/5-7/10 x2 x1 -4/10-7/10=?